OpenStudy (anonymous):

Given that one foot is approximately equal to 0.3048 meters, covert the stadium area of 3 million square feet to square meters. The stadium area is ___ meters squared.

OpenStudy (wolf1728):

One square foot = .3048² = 0.09290304 square meters That should get you off to a good start/

OpenStudy (anonymous):

then i would multiply 0.09290304 square meters by 3

OpenStudy (wolf1728):

No - more like this: 3,000,000 sq meters x 1 square foot / 0.09290304 sq meters

OpenStudy (anonymous):

its 3 million square feet though not sq meters

OpenStudy (wolf1728):

Correct but it won't be 3 million - probably closer to 30 million sq feet.

OpenStudy (anonymous):

No i was saying in the original problem

OpenStudy (anonymous):

Isn't it 3,000,000 sq feet x 1square foot /0.09290304 sq meters

OpenStudy (wolf1728):

Okay I see - the original problem states 3,000,000 sq feet. So, it would be as you said originally, it would be 3,000,000 x 0.09290304 in sq meters. ********************************************************* No - it would be 3,000,000 sq feet x 0.09290304 sq meters / 1 sq foot

OpenStudy (wolf1728):

Notice how the units cancel? (sq feet in the numerator and sq feet in the denominator).

OpenStudy (anonymous):

278,709.12 ?

OpenStudy (anonymous):

But, then rounded to nearest square meter is just 278,709

OpenStudy (wolf1728):

Yes, that is it. 278,709 Do you have other conversion problems? If so, it would be worth learning about that "units cancelling" concept.

OpenStudy (anonymous):

So 278,709 meters squared

OpenStudy (wolf1728):

or square meters

OpenStudy (anonymous):

oh just on the problem it already had M^2 i just had to fill in the number

OpenStudy (wolf1728):

okay

OpenStudy (anonymous):

I have

OpenStudy (anonymous):

Wow. Thats a lot of applications at the bottom

OpenStudy (anonymous):

Also, is the file you attached different than the one you asked at the top?

OpenStudy (wolf1728):

Quick conversion lesson: 20 inches equal how many feet? you could do this incorrectly 20 inches * 12 inches / foot The correct way is 20inches * 1 foot / 12 inches notice how the units cancel? Now for that cords of wood problem

OpenStudy (anonymous):

Yes the attachment from a minute ago is different...

OpenStudy (anonymous):

@wolf1728 You don't need to mention inches because the problem is about yards and feet?

OpenStudy (anonymous):

And you need to find out the volume of the room before you start converting into feet.

OpenStudy (anonymous):

\[V = lwh\]\[V = (4)(4)(2)\]\[V = 32\]

OpenStudy (wolf1728):

4 yards wide= 12 feet 4 yards long= 12 feet 2 yards high = 6 feet Volume = 12*12*6 = 864 cubic feet or 864 ft³ Now to convert that to cords: 864 ft³ x 1 cord / 128 ft³ = 864 / 128 cords 6.75 cords

OpenStudy (anonymous):

Ahhhhh okayyyy i seee

OpenStudy (wolf1728):

ilfy214 that inches problem was just to show how to do ANY conversion. I just posted that to show the procedure and it has nothing to do with any problem we are solving.

OpenStudy (anonymous):

I got the same as @wolf1728 but used a different method. The other way to solve the problem:\[32 \times 27 = 864ft^3\]\[\frac{ 864ft^3 }{ 128ft^3 } = 6.75cords\]

OpenStudy (anonymous):

is a square meter 1000 times as large as a square millimeter?

OpenStudy (anonymous):

and yeah @ilfy214 thats a good way also

OpenStudy (wolf1728):

No a square meter has more sq millimeters than that.

OpenStudy (anonymous):

1 square meter = 1 000 000 square millimeters

OpenStudy (anonymous):

1000000?

OpenStudy (wolf1728):

Right

OpenStudy (anonymous):

16 quarts is 15.14 liters right?

OpenStudy (wolf1728):

1 meter = 1,000 millimeters 1 square meter = 1,000 millimeters * 1,000 millimeters = 1,000,000 millimeters

OpenStudy (anonymous):

Doing lots of conversion stuff lol and i just wanna make sure i am doing it right

OpenStudy (anonymous):

One million. Why do you think it says "milli-" lolz! There is also a squared

OpenStudy (anonymous):

That was a joke... FYI

OpenStudy (wolf1728):

nslesin - remember when doing conversion problems. the units must cancel.

OpenStudy (anonymous):

Does this chart help? http://www.lessontutor.com/stairchrt3.gif

OpenStudy (anonymous):

So: A square meter is ______ times as large as a square millimeter?

OpenStudy (wolf1728):

Okay for that quarts to liters conversion 1 quart = 0.946352946 liters 16 quarts * 0.946352946 liters / 1 quart = 15.141647136 liters

OpenStudy (wolf1728):

nslesin I think you can solve that square meter problem right?

OpenStudy (anonymous):

i thought it was 1000

OpenStudy (anonymous):

If it were meter to millimeter... But these are squared

OpenStudy (anonymous):

million times greater?

OpenStudy (anonymous):

I said 1000....

OpenStudy (anonymous):

\[1 meter = 1000mm \]( milli means one thousandth) \[1meter^2 = 1000mm \times 1000mm=1,000,000 mm^2\] CORRECT

OpenStudy (anonymous):

I know sorry! I thought I said yes before

OpenStudy (anonymous):

So the answer in here is 1,000,000?

OpenStudy (anonymous):

or 1000

OpenStudy (anonymous):

@amistre64 is either of those correct?

OpenStudy (amistre64):

milli is 1000 there are 1000 mms in 1 m

OpenStudy (amistre64):



OpenStudy (anonymous):

Didn't I say that? Without a drawing? lolz

OpenStudy (amistre64):

hey, its the drawing that count :)

OpenStudy (anonymous):

True! It really is nice drawing :)

OpenStudy (anonymous):

So the answer is what exactly?

OpenStudy (amistre64):

1 square mm = .000001 sqr meters .000001 * n = 1

OpenStudy (anonymous):

where does the n come from?

OpenStudy (amistre64):

we have to scale it by something ..

OpenStudy (amistre64):

1m^2 is "n" times as large as .000001mm^2

OpenStudy (amistre64):

shoulda been m^2 on both

OpenStudy (amistre64):

1m^2 is "n" times as large as .000001m^2 .000001 * n = 1 n = 1/.000001 = 1 000 000

OpenStudy (anonymous):

Isn't the answer 1000... which is what we said before? A square meter is 1000 times as large as a square millimeter?

OpenStudy (amistre64):

no, you are trying to compare a linear measure with an area measure

OpenStudy (amistre64):

1m = 1000 mm (1m)^2 = (1000 mm)^2 1m^2 = 1000(1000) mm^2

OpenStudy (anonymous):

\[1 sqm^2 = 1,000,000 mm^2 \]

OpenStudy (amistre64):

1mm = .001 m (1mm)^2 = (.001 m)^2 1mm^2 = .000001 m^2 ergo: 1m^2 is equal to n * .000001 m^2 1/.000001 = n = 1 000 000

OpenStudy (anonymous):

so a square meter is 1,000,000 times as large has a square millimeter?

OpenStudy (anonymous):

1000^2 = 1,000,000 yup!